write the quadratic equation whose roots are 1 and 3 and whose leading coefficient is 3
Answer by Edwin McCravy(20054)


3(x-1)(x-3) = 3(x²-3x-x+3) = 3(x²-4x+3) = 3x²-12x+9

Edwin